# Ultrafast optical excitation of magnons in 2D antiferromagnetic semiconductors via spin torque mediated by unbound electron-hole pairs and excitons: Signatures in magnonic charge pumping

### v4

@@ Abstract
Recent experiments observing how femtosecond laser pulse (fsLP) excites magnons in two-dimensional (2D) antiferromagnetic (AF) semiconductors—such as CrSBr, NiPS3, and MnPS3, or their van der Waals heterostructures—suggest an important role played by excitons. However, microscopic details of such an effect remain obscure, as resonant coupling of magnons, living in the sub-meV energy range, to excitons, living in the $\sim 1$ eV range, can hardly be operative. Here, we develop a quantum transport theory of this effect, in which time-dependent nonequilibrium Green’s function (TDNEF) for electrons driven by fsLP is coupled self-consistently to the Landau-Lifshitz-Gilbert (LLG) equation describing classical dynamics of localized magnetic moments (LMMs) residing on magnetic atoms of 2D AF semiconductors. This theory explains how fsLP, of central frequency above the semiconductor gap, generates a photocurrent that becomes spin-polarized due to the background of LMMs, which, in turn, exerts spin-transfer torque (STT) onto LMMs as a genuinely nonequilibrium spintronic mechanism. The collective motion of LMMs analyzed by windowed Fast Fourier transform (FFT) decodes frequencies of excited magnons, as well as their lifetime governed by nonlocal damping with the LLG equation due to (explicitly included via TDNEGF) electronic bath. The TDNEGF part of the loop can also include excitons, which is achieved in the present study by utilizing the off-diagonal elements of nonequilibrium density matrix as the mean-field treatment of Coulomb-binding of electron-hole pairs into excitons. Finally, our theory also predicts that excited magnons will pump time-dependent charge currents into the attached electrodes, or locally within 2D AF semiconductor, thereby emitting electromagnetic radiation. The windowed FFT of these two signals contains imprints of excited magnons, as well as possible presence of excitons, which could be exploited as a novel probe in future experiments.
@@ I Introduction
The advent of two-dimensional (2D) magnetic materials [1, 2]—such as 2D antiferromagnetic (AF) semiconductors CrSBr [3, 4, 5, 6, 7, 8, 9, 10], NiPS3 [11], MnPS3 [12] or their van der Waals (vdW) heterostructures [13, 14]—has made possible recent experiments observing how femtosecond laser pulse (fsLP) excites magnons with possible exciton mediation. Excitons are quasiparticles composed of an electron-hole pair bound by Coulomb interaction [15, 16, 17, 18, 19, 20], where imaging [21, 22] shows particularly large spatial extent in the case of CrSBr. [23, 24, 25, 6] From a fundamental viewpoint, these experiments rekindle [26] interest in exciton-magnon coupling that was dormant for many decades—for example, studies [27, 28, 29, 30, 31, 32, 33] of such coupling in the 1960s and 1970s were largely motivated by absorption spectra of particular three-dimensional AF insulators like MnF2. In the contemporary context of 2D magnetic materials, coupling of magnetism and excitons has revealed novel quantum many-body effects, such as: magnon-magnon interactions [34, 35] dressed by excitons; [5] magnon-mediated exciton-exciton interactions; [10, 36] ultrafast exciton relaxation assisted by paramagnons; [20] and controlling quantum confinement (such as to monolayer) and internal structure (i.e., the strength of Coulomb interaction binding) of excitons via AF order. [24, 25] As regards applications, in magnonics [37] for classical information processing, there is a considerable effort to excite coherent magnons [38, 39] by ultrafast light, whose frequencies are as high as possible and wavelengths as short as possible [40, 12]. This is because magnons with a wavelength of $\lesssim 100$ nm would enable miniaturization of envisaged magnonic devices down to the nanoscale [41, 42]. In contrast, oscillating magnetic fields (supplied via microstrip lines or coplanar waveguides), as a standard tool to excite magnons in ferromagnets, are impractical for AF materials. Other schemes demonstrating excitation of AF magnons, such as by injecting current [43], lead to diffusive propagation of incoherent [39] (i.e., at many frequencies) magnons. Furthermore, for quantum information processing, exciton-magnon coupling offers potential for transduction [44, 45, 2] of quantum information from qubits to microwaves that would excite magnons, and then from them, via excitons, further transduction to optical photons. In turn, photons can transfer quantum information over long distances via optical fibers [45].
@@
Figure 1: Schematic view of a two-terminal setup for TDNEGF+LLG+EX [Fig. 2] calculations of ultrafast dynamics of photoexcited electrons by above CB-VB gap femtosecond laser pulse (fsLP), additionally interacting with local magnetic moments (LMMs). The central active region (CAR) consists of two layers of a 2D AF semicoductor described by the top and bottom TB chains (as inspired by the quasi-1D structure of CrSB [46, 47, 48, 49, 50, 51]) hosting two orbitals, cc(onduction) and vv(alence), per site ii. Electron spin densities (green arrows), $\langle\hat{\mathbf{s}}_{i_{c}}(t)\rangle$ and $\langle\hat{\mathbf{s}}_{i_{v}}(t)\rangle$, interact via Kondo exchange [Eqs. (10) and (11) in the Supplemental Material (SM)] with LMMs at the same site described by classical vector (red arrow) $\mathbf{M}_{i}(t)$ obeying the LLG Eq. (3). Note that LMMs are canted by applying an external magnetic field, as also used experimentally [3, 4, 5], thereby introducing noncolinearity of LMMs between the two layers. Electrons on $c$ and $v$ orbitals at the site $i$ interact via inter-orbital local Coulomb interaction of strength $U$ [Eq. (14) in the SM], which, when turned on $U>0$, binds photoexcited electrons and holes into excitons [15, 19, 16, 17]. Both chains are attached to semi-infinite ideal NM leads modeled also by 1D TB chains but without any interactions. Via such leads, any spin or charge current pumped within CAR by fsLP, or by excited magnons [52, 53, 54, 55] persisting after fsLP ceases, is drained [Fig. 4] toward macroscopic reservoirs kept at the same Fermi energy $E_{F}$ (i.e., no bias voltage is applied between the leads). We also compute EM radiation, emitted by pumped local charge currents within the CAR, and analyze its frequency content [Fig. 5].
The possibility of precise tuning of the central frequency of fsLP around subgap electronic states (like exciton or on-site $d$-$d$ transition [56, 57]) and thereby achieving control of excited magnons (such as dependence on light polarization [57]) emphasizes the key role [58] played by photoexcited electrons as mediators of magnon excitation. Thus, electrons capable of responding fast to fsLP must be explicitly included in any microscopic theory [59]. Their inclusion displaces often invoked [60, 61, 62, 63], but via intuitive reasoning, direct coupling of light to local magnetization of gapped materials—this is typically a negligible effect when examined via first-principles calculations [64]. However, the first-principles derived Hamiltonians of exciton-magnon interactions are currently lacking [65]. Furthermore, extraction of magnon and exciton spectra from a single first-principles framework, such as GW methodology [66], applied to examples of 2D AF semiconductors has revealed their vastly different energy scales [66]. This means that direct (or resonant) exciton-magnon coupling is highly unlikely [67]. One could still try to devise a Hamiltonian where off-resonant coupling emerges, but this is rarely considered as it requires special conditions [67].
Lacking such inputs and microscopic theory based on them, the magnon excitation aspect of recent experiments has been usually explained phenomenologically [3, 4, 5, 14, 9], that is, by invoking an “impulsive perturbation” [14] of unknown origin that is introduced into the Landau-Lifshitz-Gilbert (LLG) equation [68] for LMMs viewed as classical (unit) vectors $\mathbf{M}_{i}(t)$ [Fig. 1]. The agreement between such phenomenological models and experimentally excited magnon spectra implies that magnons can be approximately treated as classical spin waves [69] describable by the LLG equation, despite challenges [70, 71, 72] that AF materials can pose, especially [70] those hosting LMMs based on spin $S=1/2$ or $S=1$, for such an equation. In the case of 2D AF semiconductors with Cr as the source of LMM, the spin value is $S=3/2$ (e.g., in CrSBr, each Cr atom carries a net magnetic dipole moment of $3.03$ Bohr magnetons $\mu_{B}$, which agrees well with $S=3/2$ for Cr3+, while S and Br atoms are also slightly magnetized with 0.27-0.27 $\mu_{B}$ and 0.08-0.08 $\mu_{B}$, respectively), so that dissipative environment can induce transition [70] from quantum to classical (i.e., LLG) dynamics. We recall that in the LLG description, magnons [69, 35] emerge as collective excitations above a magnetically ordered ground state in which $\mathbf{M}_{i}(t)$ precess around a direction specified by magnetic anisotropy and/or an externally applied magnetic field. The phase of precession $\mathbf{M}_{i}(t)$ of adjacent vectors varies harmonically in space over the magnon wavelength $\lambda_{m}$. However, microscopic understanding is lacking regarding what provides the initial “kick” for LLG dynamics to be initiated. Furthermore, excited magnons could generate additional charge currents [53, 54, 52, 55] carrying their signatures within their spectral content (as initially generated photocurrent has spectral content related to much higher frequency of light [73, 74]), but their properties and usage as novel experimental probes remain unexplored.
@@
Figure 2: Flowchart of TDNEGF+LLG+EX self-consistent loop combining TDNEGF [75, 76, 77] (green box) computation [Eq. (4)] of time-dependent nonequilibrium density matrix ${\bm{\rho}}^{\rm neq}(t)$ with LLG equation [68, 69] updating (orange box) LMMs $\mathbf{M}_{i}(t)$. For this study, to the previously developed [78, 79] TDNEGF+LLG scheme, we add a computation (magenta box) employing [19] all off-diagonal elements of ${\bm{\rho}}^{\rm neq}(t)$ to describe the binding of photoexcited conduction-band electrons and valence-band holes into exctions. The loop employs time step $\delta t=0.1$ fs in both quantum (as required for numerical stability of TDNEGF calculations [75, 76, 78]) and classical LLG calculations. After each time step, we obtain time-dependent observables listed in the yellow box. In particular, STT is constructed from the expectation value of electron spin $\langle\hat{\mathbf{s}}_{i}(t)\rangle$ [Eq. (2)] and $\mathbf{M}_{i}(t)$ via Eq. (1).
@@
In this Letter, we also rely on the LLG equation, but we introduce the “kick” microscopically by self-consistently coupling [Fig. 2] the LLG equation to time-dependent quantum transport [75, 76] of photoexcited electrons, as described by the time-dependent nonequilibrium Green’s functions (TDNEGF) formalism [75, 76, 77]. Thus, our TDNEGF+LLG framework [78, 79, 52, 80, 81], or TDNEGF+LLG+EX when excitons are included [Fig. 2], makes it possible to rigorously introduce into the LLG equation the effects brought by electronic photocurrent [82]. For example, photocurrent will become spin-polarized as it propagates through the magnetically ordered background of LMMs—in addition, since LMMs within one layer of AF CAR in Fig. 1 will be spin-polarized in a direction that is noncollinear to LMMs in the second layer, electrons hopping from one layer will exert spin-transfer torque (STT) [83, 84, 85]
$\mathbf{T}_{i}(t)=J_{K}\left[\sum_{a=c,v}\langle\hat{\mathbf{s}}_{ia}\rangle(t)\right]\times\mathbf{M}_{i}(t),$ (1)
onto the LMMs of the other layer. Here, the nonequilibirum spin density is given by
$\langle\hat{\mathbf{s}}_{ia}\rangle(t)=\mathrm{Tr}_{\mathrm{spin}}[{\bm{\rho}}^{\rm neq}(t){\bm{\sigma}}],$ (2)
@@
$J_{K}$ is the Kondo exchange interaction between the spin of flowing electrons and LMMs [86, 87]; $a=c,v$ labels each of two orbitals per site [Fig. 1]; ${\bm{\sigma}}=(\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z})$ is the vector of the Pauli matrices; ${\bm{\rho}}^{\rm neq}(t)$ is the time-dependent nonequilibrium density matrix [75, 76, 88]; and the trace $\mathrm{Tr}_{\mathrm{spin}}[\ldots]$ is over quantum states in the spin space only. Importantly, without canting LMMs in Fig. 1 by the applied magnetic field, as an essential ingredient in experiments on CrSBr [3, 4, 5], STT would be zero (or unobservably small, if we include noncolinearity due to thermal fluctuations [89, 73]), so no magnons would be detectable as reported in experiments [3, 4, 5]. Such microscopically computed STT is then sent into the LLG equation

### v5

@@ Abstract
Recent experiments observing how femtosecond laser pulse (fsLP) excites magnons in two-dimensional (2D) antiferromagnetic (AF) semiconductors—such as CrSBr, NiPS3, and MnPS3, or their van der Waals heterostructures—suggest an important role played by excitons. However, microscopic details of such an effect remain obscure, as resonant coupling of magnons, living in the sub-meV energy range, to excitons, living in the $\sim 1$ eV range, can hardly be operative. Here, we develop a quantum transport theory of this effect, in which the time-dependent nonequilibrium Green’s function (TDNEGF) for electrons driven by fsLP is coupled self-consistently to the Landau-Lifshitz-Gilbert (LLG) equation describing classical dynamics of localized magnetic moments (LMMs) residing on magnetic atoms of 2D AF semiconductors. This theory explains how fsLP, of central frequency above the semiconductor gap, generates a photocurrent that becomes spin-polarized due to the background of LMMs, which, in turn, exerts spin-transfer torque (STT) onto LMMs as a genuinely nonequilibrium spintronic mechanism. By analyzing the collective motion of LMMs via the windowed Fast Fourier transform (FFT), we extract frequencies of excited magnons, as well as their lifetime governed by nonlocal damping with the LLG equation due to (explicitly included via TDNEGF) electronic bath. The TDNEGF part of the loop can also include excitons, which is achieved in the present study by utilizing the off-diagonal elements of the nonequilibrium density matrix as the mean-field treatment of Coulomb-binding of electron-hole pairs into excitons. Finally, our theory makes it possible to predict that excited magnons will pump time-dependent charge currents into the attached electrodes, or locally within the 2D AF semiconductor, thereby emitting electromagnetic radiation. The windowed FFT of these two signals contains imprints of excited magnons, as well as the possible presence of excitons, which could be exploited as a novel probe in future experiments.
@@ I Introduction
The advent of two-dimensional (2D) magnetic materials [1, 2]—such as 2D antiferromagnetic (AF) semiconductors CrSBr [3, 4, 5, 6, 7, 8, 9, 10], NiPS3 [11], MnPS3 [12] or their van der Waals (vdW) heterostructures [13, 14]—has made possible recent experiments observing how femtosecond laser pulse (fsLP) excites magnons with possible exciton mediation. Excitons are quasiparticles composed of an electron-hole pair bound by Coulomb interaction [15, 16, 17, 18, 19, 20], where imaging [21, 22] shows particularly large spatial extent in the case of CrSBr [23, 24, 25, 6]. From a fundamental viewpoint, these experiments rekindle [26] interest in exciton-magnon coupling that was dormant for many decades—for example, studies [27, 28, 29, 30, 31, 32, 33] of such coupling in the 1960s and 1970s were largely motivated by absorption spectra of particular three-dimensional AF insulators like MnF2. In the contemporary context of 2D magnetic materials, coupling of magnetism and excitons has revealed novel quantum many-body effects, such as: magnon-magnon interactions [34, 35] dressed by excitons [5]; magnon-mediated exciton-exciton interactions [10, 36]; ultrafast exciton relaxation assisted by paramagnons [20]; and controlling quantum confinement (such as to a monolayer) or internal structure (i.e., the strength of Coulomb interaction binding) of excitons via AF order [24, 25].
As regards applications, in magnonics [37] for classical information processing, there is a considerable effort to excite coherent magnons [38, 39] by ultrafast light, whose frequencies are as high as possible and wavelengths as short as possible [40, 12]. This is because magnons with a wavelength of $\lesssim 100$ nm would enable the miniaturization of envisaged magnonic devices down to the nanoscale [41, 42]. In contrast, oscillating magnetic fields (supplied via microstrip lines or coplanar waveguides), as a standard tool to excite magnons in ferromagnets, are impractical for AF materials. Other schemes demonstrating excitation of AF magnons, such as by injecting current [43], lead to diffusive propagation of incoherent [39] (i.e., at many frequencies) magnons. Furthermore, for quantum information processing, exciton-magnon coupling offers potential for transduction [44, 45, 2] of quantum information from qubits to microwaves that would excite magnons, and then from them, via excitons, further transduction to optical photons. In turn, photons can transfer quantum information over long distances via optical fibers [45].
@@
Figure 1: Schematic view of a two-terminal setup for TDNEGF+LLG+EX [Fig. 2] calculations of ultrafast dynamics of photoexcited electrons by above CB-VB gap fsLP within 2D AF semiconductor. The central active region (CAR) consists of two monolayers described by the top and bottom TB chains, as inspired by the quasi-1D structure of CrSBr [46, 47, 48, 49, 50, 51], hosting two orbitals, cc(onduction) and vv(alence), per site ii. Electron spin densities (green arrows), $\langle\hat{\mathbf{s}}_{i_{c}}\rangle(t)$ and $\langle\hat{\mathbf{s}}_{i_{v}}\rangle(t)$, interact via Kondo exchange [Eqs. (5) and (6)] with LMM at the same site described by the classical vector (red arrow) $\mathbf{M}_{i}(t)$ obeying the LLG Eq. (4). Note that LMMs are canted by applying an external magnetic field, as also used experimentally [3, 4, 5], thereby introducing noncollinearity of LMMs between the two layers. Electrons on $c$ and $v$ orbitals at the site $i$ interact via inter-orbital local Coulomb interaction of strength $U$ [Eq. (9)], which, when turned on $U>0$, binds photoexcited electrons and holes into excitons [15, 19, 16, 17]. Both chains are attached to semi-infinite ideal NM leads modeled also by 1D TB chains but without any interactions. Via such leads, any spin or charge current pumped within CAR by fsLP, or by excited magnons [52, 53, 54, 55] persisting after fsLP ceases, is drained [Fig. 4] toward macroscopic reservoirs kept at the same Fermi energy $E_{F}$ (i.e., no bias voltage is applied between the leads). We also compute EM radiation, emitted by pumped local charge currents within the CAR, and analyze its frequency content [Fig. 5].
@@
Figure 2: Flowchart of TDNEGF+LLG+EX self-consistent loop combining TDNEGF [56, 57, 58] (green box) computation [Eq. (11)] of time-dependent nonequilibrium density matrix ${\bm{\rho}}^{\rm neq}(t)$ with the LLG equation [59, 60] updating (orange box) LMMs $\mathbf{M}_{i}(t)$. For this study, to the previously developed [61, 62] TDNEGF+LLG scheme, we add a computation (magenta box) employing [19] all off-diagonal elements of ${\bm{\rho}}^{\rm neq}(t)$ to describe the binding of photoexcited conduction-band electrons and valence-band holes into excitons. The loop employs time step $\delta t=0.1$ fs in both quantum (as required for numerical stability of TDNEGF calculations [56, 57, 61]) and classical LLG calculations. After each time step, we obtain time-dependent observables listed in the yellow box. In particular, STT is constructed from the expectation value of electron spin $\langle\hat{\mathbf{s}}_{i}\rangle(t)$ [Eq. (3)] and $\mathbf{M}_{i}(t)$ via Eq. (1).
The possibility of precise tuning of the central frequency of fsLP around subgap electronic states (like exciton or on-site $d$-$d$ transition [63, 64]) and thereby achieving control of excited magnons (such as dependence on light polarization [64]) emphasizes the key role [65, 66] played by photoexcited electrons as mediators of magnon excitation. Thus, electrons capable of responding fast to fsLP must be explicitly included in any microscopic theory [67]. Their inclusion displaces often invoked [68, 69, 70, 71], via intuitive reasoning, direct coupling of light to local magnetization of gapped materials—this is typically a negligible effect when examined via first-principles calculations [72]. However, the first-principles-derived Hamiltonians of exciton-magnon interactions are currently lacking [73]. Furthermore, extraction of magnon and exciton spectra from a single first-principles framework, such as GW methodology [74], applied to examples of 2D AF semiconductors has revealed their vastly different energy scales [74]. This means that direct (or resonant) exciton-magnon coupling is highly unlikely [75]. One could still try to devise a Hamiltonian where off-resonant coupling emerges, but this is rarely considered as it requires special conditions [75].
Lacking such inputs and microscopic theory based on them, the magnon excitation aspect of recent experiments has usually been explained phenomenologically [3, 4, 5, 14, 9], that is, by invoking an “impulsive perturbation” [14] of unknown origin that is introduced into the Landau-Lifshitz-Gilbert (LLG) equation [59] for localized magnetic moments (LMMs) viewed as classical (unit) vectors $\mathbf{M}_{i}(t)$ [Fig. 1]. The agreement between such phenomenological models and experimentally excited magnon spectra implies that magnons can be approximately treated as classical spin waves [60] describable by the LLG equation, despite challenges [76, 77, 78, 79] that AF materials can pose, especially [76] those hosting LMMs based on spin $S=1/2$ or $S=1$, for such an equation. In the case of 2D AF semiconductors with Cr as the source of LMM, the spin value is $S=3/2$ (e.g., in CrSBr, each Cr atom carries a net magnetic dipole moment of $3.03$ Bohr magneton $\mu_{B}$, which agrees well with $S=3/2$ for Cr3+, while S and Br atoms are also slightly magnetized with 0.27-0.27 $\mu_{B}$ and 0.08-0.08 $\mu_{B}$, respectively), so that a dissipative environment can induce transition [76] from quantum to classical (i.e., LLG) dynamics. We recall that in the LLG description, magnons [60, 35] emerge as collective excitations above a magnetically ordered ground state in which $\mathbf{M}_{i}(t)$ precess around a direction specified by magnetic anisotropy and/or an externally applied magnetic field. The phase of precession $\mathbf{M}_{i}(t)$ of adjacent vectors varies harmonically in space over the magnon wavelength $\lambda_{m}$. However, microscopic understanding is lacking regarding what provides the initial “kick” for LLG dynamics to be initiated. Furthermore, excited magnons could generate additional charge currents [53, 54, 52, 55] carrying their signatures within spectral content (note that initially generated photocurrent has spectral content related to a much higher frequency of light [80, 81]). The properties and usage of such signatures as potential novel experimental probes remain largely unexplored.
@@
This study also employs the LLG equation, but we introduce the “kick” microscopically by self-consistently coupling [Fig. 2] the LLG equation to time-dependent quantum transport [56, 57] of photoexcited electrons, as described by the time-dependent nonequilibrium Green’s functions (TDNEGF) formalism [56, 57, 58]. Thus, our TDNEGF+LLG framework [61, 62, 52, 82, 83], or TDNEGF+LLG+EX when excitons are included [Fig. 2], makes it possible to rigorously introduce into the LLG equation the effects brought by electronic photocurrent [84]. For example, photocurrent will become spin-polarized as it propagates through the magnetically ordered background of LMMs. In addition, since LMMs within one layer of AF CAR in Fig. 1 will be spin-polarized in a direction that is noncollinear to LMMs in the second layer, electrons hopping from one layer will exert spin-transfer torque (STT) [85, 86, 87] onto the LMMs of the other layer:
$\mathbf{T}_{i}(t)=J_{K}\langle\hat{\mathbf{s}}_{i}\rangle(t)\times\mathbf{M}_{i}(t),$ (1)
where the nonequilibrium expectation value of the spin operator of an electron on the site $i$ is given by
$\langle\hat{\mathbf{s}}_{i}\rangle(t)=\left[\sum_{a=c,v}\langle\hat{\mathbf{s}}_{ia}\rangle(t)\right],$ (2)
considering that each site hosts two orbitals. Here, the nonequilibrium spin densities summed over are given by
$\langle\hat{\mathbf{s}}_{ia}\rangle(t)=\mathrm{Tr}_{\mathrm{spin}}[{\bm{\rho}}^{\rm neq}(t)\hat{\mathbf{s}}_{ia}],$ (3)
@@
$J_{K}$ is the Kondo exchange interaction between the spin of flowing electrons and LMMs [88, 89]; $a=c,v$ labels each of two orbitals per site [Fig. 1] with $\hat{\mathbf{s}}_{ia}$ being the spin operator of an electron on the site $i$ in orbital $a$; ${\bm{\rho}}^{\rm neq}(t)$ is the time-dependent nonequilibrium density matrix [56, 57, 90]; and the trace $\mathrm{Tr}_{\mathrm{spin}}[\ldots]$ is over quantum states in the spin space only. Importantly, without canting LMMs in Fig. 1 by the applied magnetic field, as an essential ingredient in experiments on CrSBr [3, 4, 5], STT would be zero (or unobservably small if we include noncolinearity only due to inevitable thermal fluctuations [91, 80]), so no magnons would be detectable as reported in experiments [3, 4, 5]. Such microscopically computed STT is then sent into the LLG equation